# Quasiparticle cooling algorithms for quantum many-body state preparation

### v1

@@ Abstract
Probing correlated states of many-body systems is one of the central tasks for quantum simulators and processors. A promising approach to state preparation is to realize desired correlated states as steady states of engineered dissipative evolution. A recent experiment with a Google superconducting quantum processor [X. Mi et al., Science 383, 1332 (2024)] demonstrated a cooling algorithm utilizing auxiliary degrees of freedom that are periodically reset to remove quasiparticles from the system, thereby driving it towards the ground state. We develop a kinetic theory framework to describe quasiparticle cooling dynamics, and employ it to compare the efficiency of different cooling algorithms. In particular, we introduce a protocol where coupling to auxiliaries is modulated in time to minimize heating processes, and demonstrate that it allows a high-fidelity preparation of ground states in different quantum phases. We verify the validity of the kinetic theory description by an extensive comparison with numerical simulations of a 1d transverse-field Ising model using a solvable model and tensor-network techniques. Further, the effect of noise, which limits efficiency of variational quantum algorithms in near-term quantum processors, can be naturally described within the kinetic theory. We investigate the steady state quasiparticle population as a function of noise strength, and establish maximum noise values for achieving high-fidelity ground states. This work establishes quasiparticle cooling algorithms as a practical, robust method for many-body state preparation on near-term quantum processors.
@@ I Introduction
Two important classes of low-energy state-preparation protocols employed in experiments with noisy intermediate-scale quantum (NISQ) devices [13] are: (i) adiabatic protocols [9, 14], which rely on the adiabatic theorem and a slow variation of Hamiltonian parameters, and have been widely used in analogue quantum simulators such as ultracold atoms [15, 16] and Rydberg atom arrays [17, 18]; (ii) variational methods [19, 20, 21], suitable for digital quantum simulators, in which the target state is prepared by applying a variational quantum circuit to a simple initial state (e.g. a product state).
@@
The unitary state-preparation protocols on NISQ processors are hindered by the presence of intrinsic decoherence. Such protocols have no native mechanism to remove excitations, so decoherence will eventually thermalize the system to a trivial state, e.g. a fully mixed state. This restricts the size of feasible circuits and limits the duration for which a quantum state can be stored on the device [22, 23]. Further, using variational circuits to prepare complex quantum states may be challenging due to flat optimization landscapes (‘barren plateaus’) [24, 25, 26].
@@
More recently, in an experiment [38] on the digital Google Quantum Processor, a physics-inspired cooling algorithm was demonstrated for cooling 1d and 2d quantum spin systems. In this experiment, the system evolved under a periodic (Floquet) quantum circuit specified by a unitary Floquet operator $\hat{U}_{S}$, constructed as a shallow quantum circuit of gates with finite, but small rotation angles. In this regime, the time evolution of local observables is well-approximated by a prethermal Hamiltonian $H_{\rm eff}$ [39]. The approximate conservation of $H_{\rm eff}$ stems from the suppression of heating processes, which are exponentially slow in the inverse rotation angles of gates in $\hat{U}_{S}$. The key idea of the implemented algorithm (displayed schematically in Fig. 1a,b — see also  [40, 41, 42, 43, 44] for related works) was to use a subset of qubits as auxiliaries, tuning their quasienergies and couplings to system qubits in order to resonantly transfer excitations from the system. The entropy of the system is removed by periodically resetting the auxiliaries. In the experiment, it was shown that the system reached a steady low-energy correlated state. Furthermore, the achieved steady states could be efficiently purified to yield high fidelities with respect to the true ground states, and it was argued that, asymptotically, the dissipative preparation is advantageous compared to variational unitary protocols.
@@
This progress calls for the development of practical, robust, dissipative many-body state preparation algorithms. In this work, we develop a theory of quasiparticle cooling on quantum processors, concentrating on efficient preparation of phases of matter with low-energy quasiparticle excitations, in realistic noisy environments. In such many-body systems, excited states at sufficiently low energy density can be described in terms of occupation numbers of non-interacting bosonic or fermionic degrees of freedom [45]. The quasiparticles generally have a finite lifetime, due to the decay to other states; however, provided this lifetime is sufficiently long, quasiparticle description can be employed. A celebrated example is the Landau Fermi-liquid theory, which provides a description of interacting electrons in metals in terms of fermionic quasiparticles [45]. We note that quasiparticle description applies to a variety of phases of matter including Fermi liquids, superconductors, and magnetically ordered states.
@@
Focusing on 1d qubit – or, equivalently, spin-$1/2$ – systems, we consider generalised auxiliary-reset algorithms encompassing the one implemented in the experiment of Ref. [38]. We derive a dissipative kinetic equation describing the quasiparticle population dynamics under the action of the protocol, which allows for a transparent physical picture on how different parameters affect the state preparation fidelity. In particular, we investigate the effects of: (i) auxiliary layout as shown in Fig. 1c,d, with auxiliaries only at the system boundaries (edge cooling setup) or with a finite density of auxiliaries in the thermodynamic limit (bulk cooling setup); and (ii) the time dependence of the system-auxiliary couplings. We find that the latter is crucial in reaching high fidelities, and we introduce a protocol with modulated couplings to drastically reduce unwanted heating processes arising due to the finite duration of a cooling cycle. Below, we characterize the protocol performance by log-fidelity per qubit, $-\log{\mathcal{F}}/N_{S}$, where $\mathcal{F}$ is the fidelity of the steady state with respect to the true ground state, and $N_{S}$ is the number of system qubits. We argue that, if the steady state is sufficiently close to the ground sate, this quantity is related to the steady-state quasiparticle density $n$ via $-\log{\mathcal{F}}/N_{S}\approx n$. Within our kinetic equation approach, we find $n\sim\exp(-\Omega(\Delta T))$ under the modulated coupling protocol, where $\Delta$ is the quasiparticle excitation gap and $T$ is the number of unitary layers before the auxiliaries are reset (see Fig. 1b). Thus, high fidelity can be achieved by increasing parameter $T$.
@@
We assess the validity of our theory first with the example of an integrable 1d transverse-field Ising model (TFIM) spin chain, where quasiparticles are infinitely long-lived and the relevant operators are known explicitly [46, 47]. This model exhibits two phases, a paramagnetic (PM) and an antiferromagnetic (AFM) phase. The distinct nature of the quasiparticle type in the two phases (magnons in the paramagnet and domain walls in the antiferromagnet), as well as spontaneous symmetry breaking in the AFM phase, lead to qualitative differences in the importance of different cooling processes. By performing extensive numerical simulations, we show that the kinetic equation correctly captures the system’s evolution and its steady state even at a relatively strong auxiliary-system coupling — the regime relevant for current experiments. We compare the experimental protocol of Ref. [38] (where the couplings were kept constant) to our modulated coupling protocol, finding that the latter allows for much higher fidelities in the limit of large $T$.
@@
Noise, inevitably present in today’s quantum simulators and processors, severely limits the performance of quantum algorithms. For example, superconducting qubits are subject to dephasing and decay processes, which introduce unwanted quasiparticles to the system. We modify the kinetic equation by incorporating noise-induced heating on top of the cooling protocol, and investigate how the non-equilibrium distribution of quasiparticle level occupations in the steady state is affected by noise. For the Floquet TFIM example, we observe that the density $n$ of quasiparticles in the steady state scales differently with the Markovian noise strength in the two phases. Characterizing decoherennce by a dimensionless parameter $\gamma$, which reflects the noise-induced error probability per qubit over one period of Floquet evolution, we find scaling $n\propto\gamma$ in the PM phase, while in the AFM phase, we observe a crossover between linear behaviour and a scaling $n\propto\sqrt{\gamma}$. This scaling behaviour was predicted in Ref. [43], and can be understood within the kinetic framework from the nature of different quasiparticle types in the two phases. Achieving low quasiparticle densities in the AFM phase therefore requires significantly weaker noise, compared to the PM phase. Nevertheless, we find that the modulated coupling protocol achieves high fidelities when the noise rate is small compared to the auxiliary-system coupling strength.
While we largely focus on cooling in the integrable TFIM model, we expect the quasiparticle cooling algorithm to be effective also for non-integrable systems. Indeed, many such systems have low-energy quasiparticles which become weakly interacting at low energy densities. Thus, at least at low energies (the most relevant regime for ground state preparation), our analysis should apply. We illustrate the versatility of our approach by studying cooling in a TFIM with broken integrability [48], where we observe similar efficiency of our protocol as in the integrable case.
@@
The rest of the paper is organized as follows: in Section II we describe the quasiparticle cooling protocol, and derive an equation that describes the evolution of the quasiparticle level populations, assuming a weak auxiliary-system coupling. We introduce the modulated coupling protocol (MCP) to suppress unwanted heating processes. In Section III, we turn to cooling the integrable TFIM. We analyse cooling with boundary auxiliaries and with a finite density of auxiliaries, for different cooling protocols. We find that the MCP leads to effective cooling, which can be accurately captured by the kinetic equation of Sec. II. Noise is introduced in Section IV: we examine how noise processes (dephasing and decay) affect the cooling, and show that these effects can be included into the kinetic description without complication. In Section V, we apply our protocol to study a non-integrable model, the TFIM with an additional longitudinal field. Finally, in Section VI, we provide concluding remarks and directions for future work.
@@ II.1 Setup and protocol
We consider the following general setup for a system of qubits, or spins-$1/2$, a particular version of which was realized in the recent experiment with the Google quantum processor [38]. The central idea follows Ref. [27], and the main developments we present here are (i) the derivation of a simplified kinetic theory, based on the quasiparticle picture, and (ii) the introduction of time-dependent system-bath coupling, which, as we will see below, is crucial to reach high fidelities.
@@
    Auxiliary qubits are initialized in a state described by a density matrix $\hat{\rho}_{B}^{0}$.
@@
    $T$ layers of unitary evolution are applied. A unitary applied at time step $1\leq\tau\leq T$ is a combination of the Floquet operator $\hat{U}_{S}$ acting on system qubits, a unitary $\hat{U}_{B}$ acting on the bath qubits, and an interaction term $\hat{U}_{\theta,\tau}$ which couples system and bath qubits. The interaction terms is generally time-dependent, and $\theta$ is the coupling strength (see Eq. (8) below).
@@
We will assume the steady state of the dynamics is unique, and is therefore reached independent of the initial state. In the analysis below we therefore take the initial state to be the maximally mixed state, $\hat{\rho}_{S}^{0}\propto\hat{I}$, unless specified otherwise.
@@
so that each auxiliary qubit simply acquires a phase in the eigenbasis of the $\hat{Z}$ operator. More complicated (e.g. interacting) baths can be considered with possible advantages. Parameter $\pi h$ is referred to as the auxiliary quasienergy and does not depend on the unitary step $\tau$.
@@
where operators $\hat{A}$ ($\hat{B}$) acts on system (bath). Note that index $a$ does not have to be identical to the index numbering auxiliary qubits.
A particular form of interaction that we will consider below arises when an auxiliary $\mathcal{A}_{j}$ is coupled to a single qubit $\mathcal{Q}_{j}$ of the system by an operator $\hat{V}$ that can generate a partial iSWAP gate:
where $\mathcal{Q}_{j}$ ($\mathcal{A}_{j}$) represent system (auxiliary) position associated with $j$-th coupling. Here $\hat{\sigma}^{\pm}=\frac{1}{2}(\hat{X}\pm i\hat{Y})$ are spin raising/lowering operators.
@@ II.2 Weak-coupling analysis
In the remaining theoretical analysis, we assume that the excitations of the many-body system are well-defined, long-lived quasiparticles, and that we are in the limit of weak system-bath coupling, $\theta|V_{j}|\ll 1$. The quasiparticle assumption holds exactly in the example of the TFIM, Eq. (5); we note, however, that interacting systems away from criticality typically have quasiparticle excitations – examples include Fermi-liquids, superconductors, and symmetry-broken magnetic phases [45]. We expect the quasiparticle cooling algorithm to remain effective for such interacting systems. Our strategy will be to derive an equation describing evolution of the quasiparticle level occupations in the weak-coupling limit.
@@
$n_{k}^{\prime}={\rm Tr}_{S}(\Phi(\hat{\rho}_{S})\hat{n}_{k})=\text{Tr}_{SB}\Big{(}\hat{\rho}_{S}\otimes\hat{\rho}_{B}^{0}\Big{[}\mathcal{\hat{U}}^{\dagger}_{T}\hat{n}_{k}\mathcal{\hat{U}}_{T}\Big{]}\Big{)},$ (14)
@@
$\Gamma_{ab}(\tau)=\text{Tr}_{B}\Big{(}\hat{\rho}_{B}^{0}\hat{B}^{a\dagger}_{I\tau}\hat{B}^{b}_{I0}\Big{)},$ (17)
and used relations $\Gamma_{ab}(\tau)=\Gamma_{ba}^{*}(-\tau)$. The second term in the r.-h.s. of Eq. (16) can be viewed as arising from the “Lamb shift” renormalization:

### v2

@@ Abstract
Probing correlated states of many-body systems is one of the central tasks for quantum simulators and processors. A promising approach to state preparation is to realize desired correlated states as steady states of engineered dissipative evolution. A recent experiment with a Google superconducting quantum processor [X. Mi et al., Science 383, 1332 (2024)] demonstrated a cooling algorithm utilizing auxiliary degrees of freedom that are periodically reset to remove quasiparticles from the system, thereby driving it towards its ground state. In this work, we develop a kinetic theory framework to describe quasiparticle cooling dynamics, and employ it to compare the efficiency of different cooling algorithms. In particular, we introduce a protocol where coupling to auxiliaries is modulated in time to minimize heating processes, and demonstrate that it allows a high-fidelity preparation of ground states in different quantum phases. We verify the validity of the kinetic theory description by an extensive comparison with numerical simulations for the examples of a 1d transverse-field Ising model, the transverse-field Ising model with an additional integrability-breaking field, and a non-integrable antiferromagnetic Heisenberg spin ladder. In all cases we are able to efficiently cool into the many-body ground state. The effects of noise, which limits efficiency of variational quantum algorithms in near-term quantum processors, are investigated through the lens of the kinetic theory: we show how the steady state quasiparticle populations depend on the noise rate, and we establish maximum noise values for achieving high-fidelity ground states. This work establishes quasiparticle cooling algorithms as a practical, robust method for many-body state preparation on near-term quantum processors.
@@ I Introduction
Two important classes of low-energy state-preparation protocols employed in experiments with noisy intermediate-scale quantum (NISQ) [13] devices are: (i) adiabatic protocols [9, 14], which rely on the adiabatic theorem and a slow variation of Hamiltonian parameters, and have been widely used in analogue quantum simulators such as ultracold atoms [15, 16] and Rydberg atom arrays [17, 18]; (ii) variational methods [19, 20, 21], suitable for digital quantum simulators, in which the target state is prepared by applying a variational quantum circuit to a simple initial state (e.g. a product state).
@@
Unfortunately, unitary state-preparation protocols on NISQ processors are hindered by the presence of intrinsic decoherence. For example, superconducting qubits are subject to dephasing and decay processes, which introduce unwanted excitations to the system. Such protocols have no native mechanism to remove these excitations, so decoherence will eventually thermalize the system to a trivial state, e.g. a fully mixed state. This restricts the size of feasible circuits and limits the duration for which a quantum state can be stored on the device [22, 23]. Further, using variational circuits to prepare complex quantum states may be challenging due to flat optimization landscapes (‘barren plateaus’) [24, 25, 26].
@@
More recently, in an experiment [38] on the digital Google Quantum Processor, a physics-inspired cooling algorithm was demonstrated for cooling 1d and 2d quantum spin systems. In this experiment, the system evolved under a periodic (Floquet) quantum circuit specified by a unitary Floquet operator $\hat{U}_{S}$, constructed as a shallow quantum circuit of gates with finite, but small rotation angles. In this regime, the time evolution of local observables is well-approximated by a prethermal Hamiltonian $H_{\rm eff}$ [39]. The approximate conservation of $H_{\rm eff}$ stems from the suppression of heating processes, which are exponentially slow in the inverse rotation angles of gates in $\hat{U}_{S}$. The key idea of the implemented algorithm (displayed schematically in Fig. 1a,b — see also  [40, 41, 42, 43, 44, 45] for related works) was to use a subset of qubits as auxiliaries, tuning their quasienergies and couplings to system qubits in order to resonantly transfer excitations from the system. The entropy of the system is removed by periodically resetting the auxiliary qubits. In the experiment, it was shown that the system reached a steady low-energy correlated state. Furthermore, the achieved steady states could be efficiently purified to yield high fidelities with respect to the true ground states, and it was argued that, asymptotically, the dissipative preparation is advantageous compared to variational unitary protocols.
@@
This progress calls for the development of practical, robust, dissipative many-body state preparation algorithms. In this work, we develop a theory of quasiparticle cooling on quantum processors, concentrating on efficient preparation of phases of matter with low-energy quasiparticle excitations, in realistic noisy environments. In such many-body systems, excited states at sufficiently low energy density can be described in terms of occupation numbers of non-interacting bosonic or fermionic degrees of freedom [46]. The quasiparticles generally have a finite lifetime, due to the decay to other states; however, provided this lifetime is sufficiently long, a quasiparticle description can be employed. A celebrated example is the Landau Fermi-liquid theory, which provides a description of interacting electrons in metals in terms of fermionic quasiparticles [46]. More broadly, quasiparticle descriptions apply to a variety of phases of matter including Fermi liquids, superconductors, and magnetically ordered states.
@@
Focusing on 1d qubit – or, equivalently, spin-$1/2$ – systems, we consider generalised auxiliary-reset algorithms encompassing the one implemented in the experiment of Ref. [38]. We derive a dissipative kinetic equation describing the quasiparticle population dynamics under the action of the protocol, which allows for a transparent physical picture on how different parameters affect the state preparation fidelity. We show that, in particular, the success of the protocol relies on (i) the local cooling operators acting to remove quasiparticles from the system, (ii) the layout of auxiliary qubits, for example with auxiliaries only at the system boundaries or with a finite density of auxiliaries in the thermodynamic limit, as shown in Fig. 1c,d, and (iii) the time dependence of the system-auxiliary couplings. The first two factors ensure that the matrix elements of the cooling operators between low-lying excited states and the ground state remain finite in the thermodynamic limit; we argue that only in topologically trivial systems is it possible to reach the ground state in constant time, while requiring the cooling operators to be local, few-qubit operators. The time-dependence of the couplings meanwhile ensures that heating transitions are suppressed relative to cooling transitions, and is crucial in reaching high fidelities.
Below, we characterize the protocol performance by log-fidelity per qubit, $-\log{\mathcal{F}}/N_{S}$, where $\mathcal{F}$ is the fidelity of the steady state with respect to the true ground state, and $N_{S}$ is the number of system qubits. Within our kinetic equation approach, we find $-\log{\mathcal{F}}/N_{S}\sim\exp(-\Omega(\Delta T))$ under the modulated coupling protocol, where $\Delta$ is the excitation gap and $T$ is the number of unitary layers before the auxiliaries are reset (see Fig. 1b), which plays the role of an inverse temperature within the modulated coupling protocol. Thus, high fidelity can be achieved by increasing the unitary circuit depth.
@@
We assess the validity of our theory first with the example of an integrable 1d transverse-field Ising model (TFIM) spin chain, where quasiparticles are infinitely long-lived and the relevant operators are known explicitly [47, 48]. This model exhibits two phases, a paramagnetic (PM) and an antiferromagnetic (AFM) phase. The distinct nature of the quasiparticle types in the two phases (magnons in the paramagnet and domain walls in the antiferromagnet), as well as spontaneous symmetry breaking in the AFM phase, lead to qualitative differences in the importance of different cooling processes. By performing extensive numerical simulations, we show that the kinetic equation correctly captures the system’s evolution and its steady state even at a relatively strong auxiliary-system coupling — the regime relevant for current experiments. We compare the experimental protocol of Ref. [38] (where the couplings were kept constant) to our modulated coupling protocol, finding that the latter allows for much higher fidelities in the limit of large $T$.
As a further test for the applicability of our protocol, we study two non-integrable and strongly interacting models: the TFIM with an integrability-breaking longitudinal field [49], and a Heisenberg spin ladder in a strongly interacting regime [50]. Since many non-integrable systems have quasiparticle excitations which become weakly interacting at low energy densities, we expect our analysis to apply, at least at low energies (the most relevant regime for ground state preparation). In both cases we observe that the system converges to an energy below the excitation gap, thus demonstrating that ground state preparation is not restricted to integrable models.
@@
While our results for the different models demonstrate the theoretical applicability of our protocol, to be practically useful the algorithm must also be able to compete against the unwanted effects of noise. We modify the kinetic equation by incorporating noise-induced heating on top of the cooling protocol, and investigate how the non-equilibrium distribution of quasiparticle level occupations in the steady state is affected by noise, for the Floquet TFIM example. We observe that the density of quasiparticles in the steady state scales differently with the Markovian noise strength in the two phases, such that achieving low quasiparticle densities in the AFM phase requires significantly weaker noise. This scaling behaviour was predicted in Ref. [43], and can be understood within the kinetic framework from the nature of different quasiparticle types in the two phases. Nevertheless, we find that the modulated coupling protocol achieves high fidelities when the noise rate is small compared to the auxiliary-system coupling strength.
@@
The rest of the paper is organized as follows: in Section II we describe the quasiparticle cooling protocol, and derive an equation for the evolution of the quasiparticle level populations, assuming weak auxiliary-system coupling. We discuss how the couplings can be tailored to remove quasiparticle excitations, and introduce the modulated coupling protocol (MCP) to suppress unwanted heating processes. In Section III, we turn to cooling the integrable TFIM. We analyse cooling with boundary auxiliaries and with a finite density of auxiliaries, for different cooling protocols. We find that the MCP leads to effective cooling, which is accurately captured by the kinetic equation of Section II. Non-integrable models are studied in Section IV, for the examples of the TFIM with an additional longitudinal field and a Heisenberg spin ladder. To check the protocol’s robustness to decoherence, in Section V we generalise the kinetic theory to include noise, and examine how dephasing and decay processes affect the cooling of the TFIM. Finally, in Section VI, we provide concluding remarks and directions for future work.
@@ II.1 Setup and protocol
We consider the following general setup for a system of qubits, or spins-$1/2$, a particular version of which was realized in the recent experiment with the Google quantum processor [38]. The central idea follows Ref. [27], and the main technical developments we present here are (i) the derivation of a simplified kinetic theory, based on the quasiparticle picture, and (ii) the introduction of time-dependent system-bath coupling, which, as we will see below, is crucial to reach high fidelities.
@@
    Auxiliary (bath) qubits are initialized in a state described by a density matrix $\hat{\rho}_{B}^{0}$.
@@
    $T$ layers of unitary evolution are applied. A unitary applied at time step $1\leq\tau\leq T$ is a combination of the Floquet operator $\hat{U}_{S}$ acting on system qubits, a unitary $\hat{U}_{B}$ acting on the bath qubits, and an interaction term $\hat{U}_{\theta,\tau}$ which couples system and bath qubits. The interaction term is generally time-dependent, and $\theta$ is the coupling strength (see Eq. (8) below).
@@
We will assume the steady state of the dynamics defined by this quantum channel is unique, and is reached independent of the initial state; we will not consider a scenario of initial states with anomalously slow relaxation to the steady state. In the analysis below we therefore take the initial state to be the maximally mixed state, $\hat{\rho}_{S}^{0}\propto\hat{I}$, unless specified otherwise.
@@
so that each auxiliary qubit simply acquires a phase in the eigenbasis of the $\hat{Z}$ operator. Parameter $\pi h$ is referred to as the auxiliary quasienergy and does not depend on the unitary step $\tau$.
@@
where operators $\hat{A}$ ($\hat{B}$) acts on system (bath). We refer to the $\hat{A}^{a}$ operators as the (system) cooling operators.
In the following we will assume that the bath operators $\hat{B}$ act simply as annihilation operators on the non-interacting bath set by Eq. (7):
$\hat{U}^{\dagger}_{B}\hat{B}^{a}\hat{U}_{B}=e^{-i\pi h}\hat{B}^{a}.$ (10)
This is simple to implement in the case of Eq. (7), by choosing $\hat{B}^{a}=\hat{\sigma}^{+}_{i_{a}}$, with $i_{a}$ labelling the auxiliary qubit, where $\hat{\sigma}^{\pm}_{i}=(\hat{X}_{i}\pm\hat{Y}_{i})/2$ are the spin raising/lowering operators. While our choice of the bath parameters may seem rather simple to simulate realistic cooling, we will see below that the complexity is shifted into the coupling choice, Eq. (8).
@@ II.2 Weak-coupling analysis
In the remaining theoretical analysis, we assume that the excitations of the many-body system are well-defined, long-lived quasiparticles, and that we are in the limit of weak system-bath coupling, $\theta|V_{j}|\ll 1$. The quasiparticle assumption holds exactly in the example of the TFIM, Eq. (5); we note, however, that interacting systems away from criticality typically have quasiparticle excitations at low energies – examples include Fermi-liquids, superconductors, and symmetry-broken magnetic phases [46]. We expect the quasiparticle cooling algorithm to remain effective for such interacting systems, as we elaborate on in Subsection II.4, and numerically demonstrate in Section IV. Our strategy will be to derive an equation describing evolution of the quasiparticle level occupations in the weak-coupling limit.
@@
$n_{k}^{\prime}={\rm Tr}_{S}(\Phi(\hat{\rho}_{S})\hat{n}_{k})=\text{Tr}_{SB}\Big{(}\hat{\rho}_{S}\otimes\hat{\rho}_{B}^{0}\Big{[}\mathcal{\hat{U}}^{\dagger}_{T}\hat{n}_{k}\mathcal{\hat{U}}_{T}\Big{]}\Big{)}.$ (14)
@@
$\Gamma_{ab}(\tau)=\text{Tr}_{B}\Big{(}\hat{\rho}_{B}^{0}\hat{B}^{a}_{I\tau}\hat{B}^{b\dagger}_{I0}\Big{)},$ (17)
and used relations $\Gamma_{ab}(\tau)=\Gamma_{ba}^{*}(-\tau)$. The fact that the autocorrelator depends on a single time variable is due to the assumption that $\hat{\rho}_{B}^{0}$ is a stationary state of $\hat{U}_{B}$. For simplicity, in the following we will also assume that $\Gamma_{ab}$ is diagonal in the coupling operators, $\Gamma_{ab}(\tau)=\delta_{ab}\Gamma_{aa}(\tau)$; this will be the case in the examples we study. The second term in the r.-h.s. of Eq. (16) can be viewed as arising from the “Lamb shift” renormalization: